OpenStudy (anonymous):

If you know the diameter of a circle, which statement describes how to find the AREA?

OpenStudy (anonymous):

A) Multiply the diameter by π. B) Square the diameter and multiply by π. C) Divide the diameter in half and multiply by π. D) Divide the diameter in half, square, and multiply by π.

OpenStudy (anonymous):

\(\pi r^2\) is the usual formula and since the radius is half of the diameter, divide by 2, square and then multiply by \(\pi\)

OpenStudy (anonymous):

btw who writes "divide the diameter in half"??

OpenStudy (anonymous):

you do not "divide in half" you can a) take half, find half of it, etc b) divide by two

OpenStudy (anonymous):

Lol My Stupid Teacher , She Also Says "The Lunch"

OpenStudy (anonymous):

Lol Thanks .

OpenStudy (anonymous):

If you know the diameter of a circle you still use the formula for finding area for a circle. The formula is; \[Area _{circle}=\pi r ^{2}\] What you know; Diameter \[\pi \approx3.14\] What you want to know; Area of the circle Process; Divide the diameter by 2 to get the radius, since; \[2r=d\] When d represents diameter. Now that you know the radius you can plug what you know into the equation; \[Area _{circle}=3.14r ^{2}\] and instead of putting \[r ^{2}\] You would put the radius you found to the power of two; \[number ^{2}\] \[Area _{circle}=3.14number ^{2}\] Answer is D) Divide diameter in half, square, and multiply by pi.

OpenStudy (anonymous):

yeah because she is out to the lunch

OpenStudy (anonymous):

l4nl wow i love that you gave me so much info as to why that was the answer thank you ^.^

OpenStudy (anonymous):

Satillite LMFAOOO Yea Shes pretty weird .